A lawn mower of width 1m is used to mow a rectangular lawn of dimension 60mx40m. The mowing is done starting from the corner towards the center. How many times should the mower be moved along the garden for 3/4th of garden to be mowed?
Answer by ankor@dixie-net.com(22740)

A lawn mower of width 1m is used to mow a rectangular lawn of dimension 60mx40m. The mowing is done starting from the corner towards the center. How many times should the mower be moved along the garden for 3/4th of garden to be mowed?
:
Assume you mean 3/4 of the lawn, not the garden. (Big trouble if you mowed the garden around here.)
:
Find the area of 3/4 of the lawn
60 * 40 * 3/4 = 1800 sq/m
:
Let x = no. of rounds made on the lawn to complete 3/4 of it
:
Each round subtracts two meters from both dimensions, so we have:
(60-2x)*(40-2x) = 1800
FOIL
2400 - 120x - 80x + 4x^2 = 1800
:
2400 - 200x + 4x^2 - 1800 = 0
Arrange as a quadratic equation
4x^2 - 200x + 600 = 0
simplify, divide by 4
x^2 - 50x + 150 = 0
:
Use the quadratic formula to find x: a=1; b=-50; c=150
:
The reasonable answer I got was approx 3.2 rounds